# An efficient probabilistic hardware architecture for diffusion-like models

### v1

@@ IV Training DTMs
The EBMs used in the experiments presented in Fig. 1 were trained by applying the standard Monte-Carlo estimator for the gradients of EBMs [64] to Eq. (4), which yields
@@
This instability in MEBMs leads to unpredictable training dynamics that can be sensitive to implementation details. An example of the training dynamics for several different types of models is shown in Fig. 5 (b). The top plot displays the quality of images generated during training, while the bottom plot shows a measure of the sampler’s mixing. Image quality is measured using the FID metric, and mixing quality is measured using the normalized autocorrelation
$r_{yy}[k]=\frac{\mathbb{E}\left[\left(y[j]-\mu\right)\left(y[j+k]-\mu\right)\right]}{\mathbb{E}\left[\left(y[j]-\mu\right)^{2}\right]},$ (15)
$\mu=\mathbb{E}(y[j]),$ (16)
where $k$ is the delay time; $y[j]$ is some low dimensional projection of the sampling chain data at iteration $j$, $y[j]=f(x[j])$; and $\mathbb{E}\left[\cdot\right]$ indicates expectation values taken over independent Gibbs sampling chains. The lower plot in Fig. 5 (b) shows the autocorrelation at a delay equal to the total number of sampling iterations used to estimate the gradients during training. Generally, if $r_{yy}$ is close to 1, gradients were estimated using far-from-equilibrium samples and were likely of low quality. If it is close to zero, the samples should be close to equilibrium and produce high-quality gradient estimates. See Appendix H for further discussion.
The destabilizing effect of non-equilibrium sampling is apparent from the blue curves in Fig. 5 (b). At the beginning of training, both quality and $r_{yy}$ increase, indicating that the multimodality of the data is being imprinted on the model. Then $r_{yy}$ becomes so large that the quality of the gradient starts to decline, resulting in a plateau and, ultimately, a degradation of the model’s quality.
@@
where $s^{t-1}=(x^{t-1},z^{t-1})$ and $x^{t-1}_{i}$ indicates the $i^{\text{th}}$ of the $M$ variables in $x^{t-1}$. This term penalizes the distance between the learned conditional distribution and a factorized distribution with identical marginals and is a form of total correlation penalty [16].
@@
We use an Adaptive Correlation Penalty (ACP) to set the $\lambda_{t}$ as large as necessary to keep sampling tractable for each layer. During training, we periodically measure the autocorrelations of each learned conditional at a delay equal to the number of sampling iterations used during gradient estimation. If the autocorrelation for the $j^{\text{th}}$ layer is close to zero, $\lambda_{j}$ is decreased, and vice-versa.
Our closed-loop control of the correlation penalty strengths is crucial, allowing us to maximize the expressivity of the EBMs while maintaining stable training. The green curves in Fig. 5 (b) show an example of training dynamics under this closed-loop control policy. Model quality increases monotonically, and the autocorrelation stays small throughout training. This closed-loop control of the correlation penalty was employed during the training of most models used to produce the results in this article, including those shown in Fig. 1.
@@
However, some subtleties prevent this specific EBM topology from being scaled indefinitely. The top plot in Fig. 5 (c) shows that scaling the number of latent variables (with fixed allowed mixing time) only improves performance if the connectivity of the graph is also scaled; otherwise, performance can decrease. This dependence makes sense, as increasing the number of latent variables in this way increases the depth of the Boltzmann machine, which is known to make sampling more difficult. Beyond a certain point, increasing the model’s ability to express complex energy landscapes may render it unable to learn, given the allowed mixing time of $K\approx 1000$. This same effect is shown in the bottom plot of Fig. 5 (c), which demonstrates that larger values of $K$ are required to support wider models holding connectivity constant.
In general, it would be naive to expect that a hardware-efficient EBM topology can be scaled in isolation to model arbitrarily complex datasets. For example, there is no good reason for which a connectivity pattern that is convenient from a wire-routing perspective would also be well suited to represent the correlation structure of a complex real-world dataset.
V Conclusion: Scaling Thermodynamic Machine Learning
@@
However, we hypothesize that the correct way to scale probabilistic machine learning hardware systems is not in isolation but rather as a component in a larger hybrid thermodynamic-deterministic machine learning (HTDML) system. Such a hybrid system integrates probabilistic hardware with more traditional machine learning accelerators.
@@
The goal of HTDML is to design practical machine learning systems that minimize the energy used to achieve desired modeling fidelity on a particular task. This efficiency will be achieved through a cross-disciplinary effort that eschews the software/hardware abstraction barrier to design computers that respect physical constraints.
Mathematically, the landscape of HTDML may be summarized as
$E_{\text{tot}}(S,D,p)=E_{\text{det}}(S,D,p)+E_{\text{prob}}(S,D,p),$ (19)
where $E_{\text{tot}}$ is the total energy consumed by some machine learning system $S$ to evaluate a model of some dataset $D$ with performance $p$. $E_{\text{tot}}$ decomposes into an energy term that comes from the deterministic computer $E_{\text{det}}$ and a term that comes from the probabilistic computer $E_{\text{prob}}$.
Only the extremes of HTDML have been explored thus far. The existing body of work on machine learning has $E_{\text{tot}}=E_{\text{det}}$ and the early demonstrations in this work have $E_{\text{tot}}=E_{\text{prob}}$. Like many engineered systems, optimal solutions will be found somewhere in the middle, where the contributions from the various subsystems are nearly balanced [6, 69, 47].
System designs between the extremes represent completely unexplored territory. Many foundational problems in HTDML still need to be solved.
@@
For example, more rigorous methods of embedding data into hardware EBMs will need to be developed to go beyond the relatively simple datasets considered here. Indeed, binarization is not viable in general, and embedding into richer types of variables (such as categorical) at the probabilistic hardware level is not particularly efficient or principled.
@@
One way to solve the embedding problem is to use a small neural network to map data into the probabilistic hardware. A naive experiment demonstrating this is shown in Fig. 6. Here, we train a small neural network to embed the CIFAR-10 dataset [43] into a binary DTM. The embedding network was trained using an autoencoder loss to binarize the data, which was then used to train a DTM. The decoder of the embedding network was then trained further using a GAN objective to increase the quality of the generated images. This training procedure is described in further detail in Appendix I.
Despite the overhead of the embedding neural network, this primitive hybrid model is efficient. As shown in the figure, the generator of the traditional GAN has to be roughly 10 times larger than the decoder of our embedding network to match the performance of the hybrid model.
@@
Figure 6: Embedding data into a DTM using a neural network. Here, we show the results of using a simple embedding model in combination with a DTM. The DTM is trained to generate CIFAR-10 images and achieves performance parity with a traditional GAN using a $\sim 10\times$ smaller deterministic neural network.
The embedding procedure employed in Fig. 6 will likely be significantly improved through further study. One major flaw with our method is that the autoencoder and DTM are not jointly trained, which means that the embedding learned by the autoencoder may not be well-suited to the way information can flow in the DTM, given its limited connectivity. The problem of using a denoising chain of EBMs in latent space has been studied in the deep learning literature, and some of this work may be leveraged to solve the embedding problem discussed here [74].
The models used in this article are small compared to what could be implemented using even an early probabilistic computer based on the DTCA. Based on the size of our RNG, it can be estimated that $\sim 10^{6}$ sampling cells could be fit into a $6\times 6\;\text{µm}$ chip (see Appendix J). In contrast, the largest DTM shown in Fig. 1 would use only around 50,000 cells.
@@
Given this gap between the size of our models and the capabilities of a potential hardware device, a natural question to study is how these probabilistic models can be scaled outside the obvious approaches considered here. This scaling likely corresponds to developing architectures that fuse multiple EBMs to implement each step in the reverse process. One possible approach is to construct software-defined graphical models of EBMs that enable non-local information routing, which could alleviate some of the issues associated with a fixed and local interaction structure.
One difficulty with HTDML research is that simulating large hardware EBMs on GPUs can be a challenging task. GPUs run these EBMs much less efficiently than probabilistic computers and the sparse data structures that naturally arise when working with hardware EBMs do not mesh well with regular tensor data types. We have both short and long-term solutions to these challenges.
@@
To address these challenges in the short term, we have open-sourced a software library [24] that enables XLA-accelerated [55] simulation of hardware EBMs. This library is written in JAX [11] and automates the complex slicing operations that enable hardware EBM sampling. We also provide additional code that wraps this library to implement the specific experiments presented in this article [29]. In the longer term, the realization of large-scale probabilistic computers, such as the one proposed in this article, using advanced transistor processes [58, 71, 46] will significantly alleviate the challenges associated with HTDML research and accelerate the pace of progress.
Appendix A Denoising Diffusion Models
@@ A.3 The Diffusion Loss
A.3.1 A Monte-Carlo gradient estimator
@@
A.4 Simplification of the Energy Landscape
@@ A.5 Conditional Generation
The training procedure works the same way as before, just using this label-augmented data. We obtain the noised training images $X_{n}$ and labels $L_{n}$ by noising each entry of $X_{0}$ and $L_{0}$ resp. independently using the forward process described in subsection A.1.2. Then we train the $n$th step model $P_{\theta_{n}}\big(V^{\text{out}}_{X}=x,V^{\text{out}}_{L}=l\,|\,V^{\text{in}}_{X}=x^{\prime},V^{\text{in}}_{L}=l^{\prime}\big)$ to approximate (in terms Kullback-Leibler divergence) the distribution $\mathbb{P}\big(X_{n}=x,L_{n}=l\,|\,X_{n+1}=x^{\prime},L_{n+1}=l^{\prime}\big)$.
@@ A.6 Learning the marginal
Appendix B Hardware accelerators for EBMs
@@ B.1 Quadratic EBMs
In the main text, we discuss Boltzmann machines, which involve only binary random variables and are the most basic form of quadratic EBM. The Conditional Update for Boltzmann Machines requires biasing a Bernoulli random variable according to a sigmoid function of a linear combination of the model parameters and the binary neighbor states, as shown in the main text, Eq. 11. This conditional update is efficiently implementable using an RNG with a sigmoidal bias and resistors, as discussed in section J.
@@ B.1.1 Potts models
B.1.2 Gaussian-Bernoulli EBMs
@@
Appendix C A hardware architecture for denoising
@@
C.1 Implementation of the forward process energy function
@@
C.2 Implementation of the marginal energy function
@@
Appendix D Energetic analysis of the hardware architecture
@@ D.2 Local communication
Given the connectivity patterns shown in table  1, it is possible to estimate the total capacitance $C_{n}$ associated with the wire connecting a node to all of its neighbors,
@@
where $\ell\approx 6\mu m$ is the sampling cell side length, and $\eta\approx 350\text{aF}/\mu m$ is the wire capacitance per unit length in our process, see Fig. 11 (b). $a_{i}$ and $b_{i}$ are the $x$ and $y$ components of the $i^{th}$ connection rule, as described in section  C.C.2.

### v2

@@ IV Training DTMs
The EBMs used in the experiments presented in Fig. 1 were trained by applying the standard Monte Carlo estimator for the gradients of EBMs [62] to Eq. (4), which yields
@@
This instability in MEBMs leads to unpredictable training dynamics that can be sensitive to implementation details. An example of the training dynamics for several different types of models is shown in Fig. 5 (b). The top plot displays the quality of images generated during training, while the bottom plot shows a measure of the sampler’s mixing. Image quality is measured using the FID metric, and mixing quality is measured using normalized autocorrelation $r_{yy}$ (see Appendix G). The lower plot in Fig. 5 (b) shows the autocorrelation at a delay equal to the total number of sampling iterations used to estimate the gradients during training. Generally, if $r_{yy}$ is close to 1, gradients were estimated using far-from-equilibrium samples and were likely of low quality. If it is close to zero, the samples should be close to equilibrium and produce high-quality gradient estimates. See Appendix G for further discussion.
The destabilizing effect of non-equilibrium sampling is apparent from the blue curves in Fig. 5 (b). At the beginning of training, both quality and $r_{yy}$ increase, indicating that the multimodality of the data is being imprinted on the model. As training progresses and the energy landscape becomes more rugged, $r_{yy}$ becomes so large that the quality of the gradient starts to decline, resulting in a plateau and, ultimately, a degradation of the model’s quality.
@@
where $s^{t-1}=(x^{t-1},z^{t-1})$ and $x^{t-1}_{i}$ indicates the $i^{\text{th}}$ of the $M$ variables in $x^{t-1}$. This term penalizes the distance between the learned conditional distribution and a factorized distribution with identical marginals and is a form of total correlation penalty [15]. See Appendices H.H.2 and M for further discussion of this penalty and its relationship to mixing time and sampling effort.
@@
We use an Adaptive Correlation Penalty (ACP) to set the $\lambda_{t}$ as large as necessary to keep sampling tractable for each layer. During training, we periodically measure the autocorrelations of each learned conditional at a delay equal to the number of sampling iterations used during gradient estimation. If the autocorrelation for the $j^{\text{th}}$ layer is close to zero, $\lambda_{j}$ is decreased, and vice versa.
Our closed-loop control of the correlation penalty strengths is crucial, allowing us to maximize the expressivity of the EBMs while maintaining stable training. The green curves in Fig. 5 (b) show an example of training dynamics under this closed-loop control policy. Model quality increases monotonically, and the autocorrelation stays small throughout training. This closed-loop control of the correlation penalty was employed during the training of most models used to produce the results in this article, including those shown in Fig. 1. Further analysis of the ACP and its effect on mixing is given in Appendices H.H.2 and M.
@@
However, some subtleties prevent this specific EBM topology from being scaled indefinitely. The top plot in Fig. 5 (c) shows that scaling the number of latent variables (with fixed allowed mixing time) only improves performance if the connectivity of the graph is also scaled; otherwise, performance can decrease. This dependence makes sense, as increasing the number of latent variables in this way increases the depth of the Boltzmann machine, which is known to make sampling more difficult. Beyond a certain point, increasing the model’s ability to express complex energy landscapes may render it unable to learn, given the allowed mixing time of $K\approx 1000$. This same effect is shown in the bottom plot of Fig. 5 (c), which demonstrates that larger values of $K$ are required to support wider models while holding connectivity constant. Further study of how performance scales with sampling time $K$ and the number of denoising steps $T$ is provided in Appendix M.
In general, it would be naive to expect that a hardware-efficient EBM topology can be scaled in isolation to model arbitrarily complex datasets. For example, there is no good reason why a connectivity pattern that is convenient from a wire-routing perspective would also be well suited to represent the correlation structure of a complex real-world dataset.
V Hybrid Thermodynamic-Deterministic Machine Learning
@@
However, we hypothesize that the correct way to scale probabilistic machine learning hardware systems is not in isolation but rather as a component in a larger hybrid thermodynamic-deterministic machine learning (HTDML) system. Such a hybrid system integrates probabilistic hardware with more traditional machine learning accelerators. This hybrid approach is particularly important for moving beyond binarized grayscale datasets such as Fashion-MNIST toward richer, higher-dimensional data.
@@
The goal of HTDML is to design practical machine learning systems that minimize the energy used to achieve desired modeling fidelity on a particular task. This efficiency will be achieved through a cross-disciplinary effort that eschews the software/hardware abstraction barrier to design computers that respect physical constraints and leverage each type of hardware where it is most effective.
@@
For example, more rigorous methods of embedding data into hardware EBMs will need to be developed to go beyond the relatively simple datasets considered here. Indeed, binarization is not viable in general, and embedding into richer types of variables (such as categorical) at the probabilistic hardware level is not particularly efficient or principled. Instead, one may try a hybrid (HTDML) approach where a small neural network (the deterministic part) learns to embed the data into a binary latent space and a DTM (the thermodynamic part) can then be trained inside this binary space. A similar approach has been very beneficial in the context of diffusion models [71], to the point that now many state-of-the-art diffusion models operate in the latent space of a pre-trained VAE.
@@
We show the results of a quick initial attempt at such hybrid models in Fig. 6. Here, we train a small neural network to embed the CIFAR-10 dataset [42] into a binary DTM. The embedding network was trained using an autoencoder loss to binarize the data, which was then used to train a DTM. The decoder of the embedding network was then trained further using a GAN objective to increase the quality of the generated images. This training procedure is described in further detail in Appendix J. Example CIFAR-10 images generated by our model can be found in Appendix N.
Despite the overhead of the embedding neural network, this primitive hybrid model is efficient. As shown in the figure, to achieve equivalent performance with a purely NN-based GAN, the deterministic part of that GAN would need to have roughly ten times more parameters than the deterministic part of our hybrid model. Informally, one might say that therefore the DTM holds about 90% of the expressivity of our model while the neural network accounts for only about 10% of the parameters.
We should again stress that this is just a naive first attempt at a hybrid model, and we believe that further research into hybrid thermodynamic–deterministic models can lead to even more significant efficiency gains and scale this approach to larger tasks. For example, one major flaw of the embedding procedure employed in Fig. 6 is that the autoencoder and DTM are not jointly trained, which means that the embedding learned by the autoencoder may not be well-suited to the way information can flow in the DTM, given its limited connectivity. Finding a good way to jointly train them seems like a promising future research direction.
@@
Figure 6: Embedding data into a DTM using a neural network. Here, we show the results of using a simple embedding model in combination with a DTM. The DTM is trained to generate CIFAR-10 images and achieves performance parity with a traditional GAN using a $\sim 10\times$ smaller deterministic neural network in terms of parameter count.
VI Conclusion
In this paper, we have given an early view of how one might design generative machine learning algorithms and hardware jointly for thermodynamic probabilistic computers. At the modeling level, we introduced Denoising Thermodynamic Models (DTMs), which repurpose hardware EBMs as denoising steps rather than monolithic models of the data distribution and thereby avoid the mixing–expressivity tradeoff that limits monolithic EBMs. At the hardware level, we proposed the Denoising Thermodynamic Computer Architecture (DTCA), which implements DTMs using sparse, locally connected Boltzmann machines driven by an all-transistor RNG with a tunable bias and short autocorrelation time. By combining these ingredients with an adaptive correlation-penalty training procedure, we estimate that a DTCA device could match the performance of GPU-based generative models on binarized Fashion-MNIST while using roughly four orders of magnitude less energy per generated sample.
It is important to note that the experiments presented in this paper were done using simulations of thermodynamic hardware, running on a classical computer, which substantially limited the maximum size of the thermodynamic model we could represent. Specifically, the largest DTM shown in Fig. 1 uses only around 50,000 nodes in total, which is much smaller than what we expect the capacity of early DTCA chips will be. In fact, based on the size of our RNG, it can be estimated that $\sim 10^{6}$ sampling cells could be fit into a $6\times 6\;\text{µm}$ chip (see Appendix K). Furthermore, measurements of our RNG suggest that early DTCA chips will be about two to three orders of magnitude faster than the simulations used in this article. As shown in fig.˜17, using more Gibbs steps during training and inference as well as using more denoising steps both increase model performance (albeit with diminishing returns). Furthermore, as shown in fig.˜5 (c, bottom), as the model gets larger, increasing the number of Gibbs steps has a larger impact on performance. Realizing large-scale probabilistic computers using advanced transistor processes [56, 68, 45] will therefore both increase the capabilities of the DTM algorithm presented here and substantially speed up future research into other thermodynamic algorithms. The present work should thus be viewed as establishing a qualitative scaling picture and design template rather than as an optimized end-point.
@@ VI Conclusion
Given the gap between the size of our models and the capabilities of a potential hardware device, a natural question to study is how these probabilistic models can be scaled beyond the obvious approaches considered here. This scaling likely corresponds to developing architectures that fuse multiple EBMs to implement each step in the reverse process. One possible approach is to construct software-defined graphical models of EBMs that enable non-local information routing, which could alleviate some of the issues associated with a fixed and local interaction structure. Such research would go hand-in-hand with HTDML, since classical neural networks could be used to learn an embedding that hierarchically distributes the generative task between several EBMs in a way similar to nVAE [65] or StyleGAN [39]. Our preliminary CIFAR-10 experiments already show that pairing a relatively small deterministic front-end with a DTM back-end can offload most of the generative burden onto the thermodynamic component.
@@
To address the current unavailability of probabilistic hardware, we have open-sourced a software library [23] that enables XLA-accelerated [53] simulation of hardware EBMs. This library is written in JAX [10] and automates the complex slicing operations that enable hardware EBM sampling. We also provide additional code that wraps this library to implement the specific experiments presented in this article [28]. By making these tools available, we aim to lower the barrier for others to explore alternative thermodynamic algorithms, hardware topologies, and hybrid architectures, and to stress-test and refine the DTCA design in a broader range of applications.
In short, DTMs and the DTCA provide a concrete path toward generative models and hardware that are co-designed for thermodynamic probabilistic computation, rather than retrofitted to existing digital accelerators.
VII Declarations
VII.1 Data Availability and Code Availabilty
The code which reproduces our results and data is available at github.com/pschilliOrange/dtm-replication.
VII.2 Acknowledgements
This research was funded by Extropic Corp.
VII.3 Author Contributions
AJ and TM contributed equally, with TM doing most of the writing and energy estimation work, whereas AJ contributed most to the algorithms for training and evaluating the models presented. OL evaluated the performance of competitive models that we were comparing against. AG helped design some of the mentioned integrated circuits. PS helped with some experiments, and formatted the research code (which was published on GitHub) to be easy to understand and use. IC advised TM at the outset of this project. GV provided suggestions at the idea stage.
VII.4 Competing Interests
All authors are part of Extropic corp., which produces and aims to sell products using some of the technologies described in the paper.
Appendix A List of Acronyms
Acronym Meaning
AI Artificial Intelligence
ACP Adaptive Correlation Penalty
DTM Denoising Thermodynamic Model
DTCA Denoising Thermodynamic Computer Architecture
EBM Energy-Based Model
MEBM Monolithic Energy-Based Model
FID Fréchet Inception Distance
FLOP Floating-Point Operation
GAN Generative Adversarial Network
GPU Graphics Processing Unit
HTDML Hybrid Thermodynamic–Deterministic Machine Learning
LLM Large Language Model
NN Neural Network
PDK Process Development Kit
RNG Random-Number Generator
VAE Variational Autoencoder
DDPM Denoising Diffusion Probabilistic Model
Table 1: Acronyms used throughout this work.
Appendix B Denoising Diffusion Models
@@ B.3 The Diffusion Loss
B.3.1 A Monte-Carlo gradient estimator
@@
B.4 Simplification of the Energy Landscape
@@ B.5 Conditional Generation
The training procedure works the same way as before, just using this label-augmented data. We obtain the noised training images $X_{n}$ and labels $L_{n}$ by noising each entry of $X_{0}$ and $L_{0}$ resp. independently using the forward process described in subsection B.1.2. Then we train the $n$th step model $P_{\theta_{n}}\big(V^{\text{out}}_{X}=x,V^{\text{out}}_{L}=l\,|\,V^{\text{in}}_{X}=x^{\prime},V^{\text{in}}_{L}=l^{\prime}\big)$ to approximate (in terms Kullback-Leibler divergence) the distribution $\mathbb{P}\big(X_{n}=x,L_{n}=l\,|\,X_{n+1}=x^{\prime},L_{n+1}=l^{\prime}\big)$.
@@ B.6 Learning the marginal
Appendix C Hardware accelerators for EBMs
@@ C.1 Quadratic EBMs
In the main text, we discuss Boltzmann machines, which involve only binary random variables and are the most basic form of quadratic EBM. The Conditional Update for Boltzmann Machines requires biasing a Bernoulli random variable according to a sigmoid function of a linear combination of the model parameters and the binary neighbor states, as shown in the main text, Eq. 11. This conditional update is efficiently implementable using an RNG with a sigmoidal bias and resistors, as discussed in section K.
@@ C.1.1 Potts models
C.1.2 Gaussian-Bernoulli EBMs
@@
Appendix D A hardware architecture for denoising
@@
D.1 Implementation of the forward process energy function
@@
D.2 Implementation of the marginal energy function
@@
Appendix E Energetic analysis of the hardware architecture
@@ E.2 Local communication
Given the connectivity patterns shown in table  2, it is possible to estimate the total capacitance $C_{n}$ associated with the wire connecting a node to all of its neighbors,
@@
where $\ell\approx 6\mu m$ is the sampling cell side length, and $\eta\approx 350\text{aF}/\mu m$ is the wire capacitance per unit length in our process, see Fig. 11 (b). $a_{i}$ and $b_{i}$ are the $x$ and $y$ components of the $i^{th}$ connection rule, as described in section  D.D.2.